lemon1boy1:

Quinn is flying a kite. The angle of elevation formed by the kite string and the ground is 44°, and the kite string forms a straight segment that is 90 feet long. Explain how to find the distance between the ground and the kite. Include a description of the triangle you drew to help you solve, including the variables and measurements you assigned to each side and angle. Round your answer to the nearest foot.

snowflake0531:

 x is what we need to solve for

snowflake0531:

I hope you understand the image

snowflake0531:

\[sin(\theta)=\frac{opposite}{hypotenuse}\\cos(\theta)=\frac{adjacent}{hypotenuse}\\tan(\theta)=\frac{opposite}{adjacent}\] theta is the 44 degrees, and we have the opposite side and the hypotenuse, so we use sin so taking the first formula, sin(44) = x/90 x is sin(44) * 90, which you can solve with a calculator